# Towards the Terminator Economy: Assessing Job Exposure to AI through LLMs

### v2

@@ 3.3 Step 3: Compute the AI sustitutability
The TEAI index provides a measure of exposure to AI which does not specifically address the issue of task substitution or replacement.
To deal with this, we leverage the information contained in LLM motivations obtained in the fist step. We start by considering the three motivations provided by the three LLMs in evaluating each task. We asked another LLM to generate a single motivation that summarizes the different motivations regarding the exposure rate for each task of each occupation. This summary provides a description of the task activities along with a description of what AI can contribute to task implementation.
We provided this summary along with the task title and the job occupation title to a “master” LLM [^10] A larger model than the one used in the first phase (72B versus 7B). by asking it, based on the input information, the level of engagement for the task by the AI, on a scale from 1 which denotes “no engagement” to 5 which denotes “complete replacement of the human being in performing the task”. The model was prompted to provide a motivation for the evaluation undertaken. In addition, since many tasks are in theory completely implementable by AI, we also asked the model to return a Boolean flag equal to 0 for fully automated tasks that do not require any human supervision and equal to 1 for tasks where at least some human supervision is required. [^11] As this time we did not use different LLMs, to limit hallucinations, in accordance with the literature (Ji et al., 2023), we also asked the model to return the elements given as input, i.e., occupation title and task title. During the evaluation phase, we put controls on the returned strings. The basic idea is that if the model hallucinated itself, it should return a different occupation or task title than the ones given as input.
In this way, we obtained a specific measure of task substitutability that we aggregated at the occupation level using eq (2), obtaining an index of task replacement by AI (TRAI ).
@@ 4 Human Evaluation of TEAI/TRAI indexes
To evaluate the outcomes provided by LLMs, we conducted a user study by recruiting 12 evaluators through Prolific (Palan and Schitter, 2018), an online service offering participants for research studies. It has been recognised as a cost-effective platform that provides high-quality data for online research, making it suitable for studies where reliability and data integrity are priorities (Douglas et al., 2023), and used in several computer-science related user evaluations, see, e.g. (De Filippo et al., 2024; Koppol et al., 2021).
Setting the user study
We sampled occupations from the O*NET database by SOC first-digit, selecting 20% of occupations and choosing two tasks for each occupation: one with the highest and one with the lowest TEAI/TRAI indexes. This resulted in a total of 200 occupations and 400 tasks. The sample was further divided into four chunks of 100 tasks each. Each sample was composed of job occupation title, task title, TEAI value, TEAI natural language motivation, TRAI value and TRAI natural language motivations. [^12] To avoid potential biases in human judgment, the evaluators were not informed that AI systems had generated the indexes and summaries they reviewed Each chunk was assigned to exactly three participants, whom we asked to consider each task for each job occupation and evaluate the TEAI and TRAI , with a nominal value [^13] I disagree with both; I agree only with TEAI; I agree only with TRAI; I agree with both indicating their degree of agreement with the indices. We aggregate the results obtaining agreement performance for TEAI and TRAI . It is important to highlight that these results are not intended to represent true accuracy, but a consistency with what our methodology has estimated. Consensus among the participants was also calculated using the equation 1 to assess the extent to which participants agreed on the assignment of evaluation values. Results are shown in Tab. 3.
Results
On average, evaluators tend to agree with (i) TEAI index 75% of the time (48%+27%); (ii) 71% of the time with TRAI (48%+23%), whilst (iii) only 2% of the time they disagree with the indicators. However, the reader may notice that the level of consensus among reviewers is lower than expected. This is largely because the topic at hand — the impact of AI on tasks and professions — is inherently subjective and influenced by human variability, convictions, and expectations, making evaluations highly personal and, at times, open to interpretation. Nonetheless, the results show that our indexes are meaningful as they effectively capture the current capabilities of state-of-the-art AI technologies in supporting and replacing tasks within various jobs. As a result, TEAI , TRAI , and the natural language explanations provide a valuable framework for individuals and organizations to develop policies that integrate AI into careers and the workforce.
@@ Results
Agree with both 35.8 53.8 32.0 65.9 48 .69 .72 .54 .72 .68
Agree more with TEAI 28.0 19.0 45.5 17.0 27 .66 .57 .67 .52 .61
Agree more with TRAI 35.1 26.9 20.4 11.2 23 .75 .76 .70 .64 .71
Disagree with both 1.1 0.4 2.2 5.9 2 .50 .27 .41 .28 .36
Table 3: Human evaluation results
5 Experimental results and outcomes in the US labor market
5.1 Benchmarking evaluation
First, we compare our AI index with other existing measures in the literature. Figure 3 shows the correlation between the TEAI index and the well-known measure developed by (Frey and Osborne, 2017), the AI exposure index by (Felten et al., 2021) and by (Webb, 2023), and the offshorability index developed by (Acemoglu and Autor, 2011). The pairwise correlation is always statistically significant at 5%. It is higher for the AIOE index, much lower for the AI Webb and the offshorability index, and negative for the Frey-Osborne index. This means our measure is broadly consistent with existing measures but captures different elements of the relationship between AI and the labor market. The negative correlation with the Frey and Osborne index can be explained by the latter being a measure of exposure to robotics and computerization and more focused on routine tasks. At the same time, generative AI is more focused on cognitive/non-routine tasks.
@@ 5.2 AI and skills
Next, we explore the relationship between our TEAI index and different skills. Figure 4 displays scatterplots comparing the TEAI index with the intensity of different skill types at the occupation level derived from (Acemoglu and Autor, 2011). The graph shows the peculiar nature of AI technologies, which are positively correlated with cognitive analytical and interpersonal skills, but negatively correlated with routine manual skills and non-routine manual skills that require physical adaptability. Surprisingly, the correlation with cognitive routine skills is only weakly positive, while it is positive for non-routine manual skills requiring interpersonal adaptability.
Figure 4: Correlation with different skill intensity measures. Each dot represents a SOC occupation.
The results of the figure are purely descriptive; therefore, we add a more robust analysis by extracting from ONET the detailed skills associated with each occupation. We group the skills into 4 classes: Cognitive, Social, Problem Solving and Management, and Technical skills. We then develop a skill relevance index for each class at the occupation level by weighting each skill according to its level and importance. [^14] As provided by ONET. The skill relevance index is constructed as follows:
@@
where each observation is a SOC occupation ($i$), $TEAI_{i}$ is our measure of AI exposure, $\bm{S}$ is a 4x1 column vector of skill relevance for each of the skill classes described above at the occupation level, $\bm{\beta}$ is a 4x1 column vector of coefficients, $O$ defines occupation dummies and $\epsilon_{i}$ the error term. We saturate the model with more detailed dummies up to the fifth digit; therefore, the results are identified within each group variation. Table 4 shows the results. The TEAI index is positively correlated with cognitive, problem-solving and managerial skills, but negatively correlated with social skills, as expected. The relationship with technical skills is very weak and does not survive the inclusion of detailed SOC occupation dummies.
@@
Table 4: OLS estimates of TE-AI index on measures of skill intensity
@@ 5.3 AI employment and wages
Finally, we explore the relationship between TEAI and labor market outcomes. We start by analyzing the size and the characteristics of workers exposed to AI technologies. First, we divide the distribution of TEAI scores into three tertiles representing high, medium, and low AI exposure. We then compute the degree of exposure of the US population using BLS employment data. Finally, we distinguish between occupational groups and by skill groups within each tertile.
@@
Figure 5: Exposure to TEAI index by SOC group (Fig. 5(a)) and by skill intensity (Fig. 5(b)). US BLS employment. Values in millions of workers. Each bar represents a tertile of the TEAI score distribution.
@@
Figures 5(a) and 5(b) show the results. Overall, in 2023, 34% of US employment is highly exposed to AI technologies, while medium and low exposure represent 32% and 34%, respectively. Our findings do not suggest a polarizing effect of AI exposure as found by (Frey and Osborne, 2017); on the contrary, AI seems to have a more balanced impact on the labor market. This is because our indicator is able to capture recent advances in AI, such as LLMs, which have affected occupational groups such as management, business, administration, and finance, as well as ICT and science, which are intensive in non-routine cognitive tasks. For example, AI technologies are increasingly being used to diagnose diseases, write reports, code, or brainstorm ideas in management and business. On the contrary, previous studies that focus more on the impact of AI on routine tasks find that these tasks and occupations are less exposed to AI.
Grouping occupations by skill intensity shows that in the group with high exposure to AI, 88% of employment is in high-skill jobs; in the group with medium exposure, 53% of employment is in medium-skill jobs while 40% is in high-skill jobs. In the group with the lowest exposure to AI, 67% are medium-skill jobs and 25% low-skill jobs. Overall, AI exposure disproportionately affects high-skill jobs, which are characterized by the competencies most affected by AI technologies. Next, we analyze the relationship between AI exposure and workers’ characteristics.
Figure 6 shows that TEAI exposure is higher for workers with a high level of education, particularly graduates and postgraduates. There is a slight increase in exposure with age, although the variation in exposure is really limited after the age of 30. Men are more exposed than women at all ages.
@@
Panel a) shows coefficients of regression of education categories on TEAI exposure (in percentiles). Covariates include age and sex. Estimates control for occupation(4d), industry(3d), state and year fixed effects. ACS individual weights are used. Robust standard errors are clustered at the industry level. Panel b) is a binscatter. The x-axis is the average age of workers in an industry-occupation-state observation in the 2022-18 ACS 5 years sample. Biscatter is computed considering education as a covariate. ACS individual weights are used.
Finally, we assess the relationship between AI exposure, employment, and wages. To compute the medium-term effect of AI in a flexible way, allowing for changes during the estimation period, we compute the log change in employment and wages over a 4-year rolling window from 2003 to 2023.151515The last 4Y variation is therefore 2019-2023. Therefore, for each 4Y window, we run the following regression:
where $\Delta y_{i,j}$ denotes the 4-year change in log employment and log wages in sector $j$ for occupation $i$, $\bm{Z}$ is a column vector of controls, $\delta_{i}$ denote occupation dummies, $\eta_{j}$ sector dummies and $\epsilon_{i,j}$ the error term. To control for possible endogeneity and omitted variable problems, we include as controls the initial level of employment, the initial level of wage, and wage squared. We also include detailed sector (NAICS) and occupation (SOC) fixed effects in the regression and cluster the errors at the sector level. Figure 7 shows that exposure to AI positively correlates with employment and wage growth throughout the entire period. This suggests that AI technologies complement labor and increase productivity, thereby boosting employment and wages in occupations with greater exposure to AI.
The presence of detailed controls at the industry and occupation level allows us to control for factors on the production side (changes in output across industries), on the demand side (changes in product demand across industries) and on the labor supply side (changes in employment across industries and occupations) that are unrelated to AI technologies and that could affect wages and employment. Moreover, the focus on a relatively short period isolates our results from long-term trends within industries and occupations.
Therefore, the positive relationship between employment and wages and AI exposure should be interpreted as meaning that occupations more exposed to AI have stronger employment and wage growth within the occupation and sector. Our results contrast with those obtained by (Acemoglu et al., 2022; Webb, 2023), who find a negative relationship between employment and wages. The potential reconciliation between our findings and theirs lies, on the one hand, in our construction of a different measure of AI exposure that emphasizes more recent advances in AI. On the other hand, our analysis focuses on changes over the last 20 years, while their analysis takes a longer-term perspective, focusing on changes over several decades.
@@
This figure plots the effect of AI score on employment and wage growth. Estimates are obtained from equation 4, with rolling regression coefficients and 95% confidence intervals of 4-year windows, starting in 2003-2007 and ending in 2019-2023. The point estimate refers to TEAI score, and the dependent variables are annual percentage growth rates of employment and wages. Employment regression includes the log of the initial period of employment. Wage regression includes a log of initial period employment, log initial period wage and log initial period wage squared. All the regressions include occupation (SOC 4 digit) and sector (NAICS 3 digit) fixed effects. Robust standard errors clustered at NAICS level.
@@ 5.4 Task exposure or substitutability?
Whether AI will be an opportunity or a threat in the future will hinge on whether it complements or substitutes human labor.
@@
The measure we have developed so far is relatively agnostic in this respect, as we cannot yet disentangle substitutability from complementarity. In other words, a high TEAI exposure measure does not necessarily imply full substitution of labor by technology, which, on the contrary, may rather complement human activities, leading to higher productivity without displacing labor.
As described in section 3 the TRAI index is specifically designed to assess task replacement.
Figure 8: Task exposure and task substitutability
This figure plots the TEAI index and the task replacement index TRAI . Each dot represents an O*NET occupation. Division lines are constructed using median values of the distribution of both indexes. An interactive visualisation and the full list of occupations and AI indexes are available at https://terminatoreconomy.com/home
(a) Management, Business, Admin, Finance & ICT Science
(b) Education, Social, Legal, Art & Health & Care
(c) Other Services vs Sales
(d) Farming & Fishing & Construction & Mining
(e) Production & Repair &Transport
Figure 9: Exposure and task substitutability by occupation groups.
This figure plots the TEAI index and the replacement index TRAI . Each panel represents a pair of occupation groups. Each dot represents an occupation. Division lines are constructed using median values.
@@
What is the relationship between the TRAI and the TEAI index? On the one hand, the TRAI index can be considered as a sub-product of the TEAI index, as it is constructed by using the motivation summaries that LLMs produce in developing the TEAI . On the other hand, it differs because the task description that it analyzes is different (the TEAI uses ONET task description TRAI uses the summary of LMMs descriptions and motivations) and is designed to assess task substitutability, as the prompt asks LLM to specifically assess the possibility of replacing the human in task performance.
@@
Figure 8 displays the two indexes. Both scores have been rescaled on a scale of 0-100 for better readability. Please note that the scaling does not imply an exact percentage. For example, a score of 80 does not mean that there is an 80% exposure or a replacement rate of 80%. Also, while the figure shows different jobs, our replacement and exposure measures are constructed at task level. So, even if an occupation has a high level of replacement, it does not mean that it can be completely replaced by AI. Instead, it means that AI can perform a large number of tasks in that occupation, either completely or partially. The figure shows that the two measures are highly correlated, i.e. that high values of the TEAI index are generally associated with high values of the replacement index.
@@
Grouping occupations improves the explanatory power of the analysis. In Figure 9 we isolate different occupation groups using the same taxonomy as in Figure 5(a). For example, most ICT and science occupations display a high level of exposure and task substitutability. This means that for these occupations, AI can execute several tasks with a high degree of confidence. Human activity will still be required, and it is likely that these professionals will concentrate on more critical and high value added tasks (such as supervision, analysis, developing solutions, etc.) leaving other tasks to AI. Analogously most teachers are characterized by a high level of exposure by a low level of substitutability: they are likely to work with AI while AI will replace few tasks. Similarly, several health and care occupations are characterized by low level of exposure and substitutability with few exceptions, such as “Medical record specialists” whose tasks are highly replaceable by technology. This provides some insights about the complementarity vs substitutability debate. Complementarity is often interpreted as AI and humans sharing the same activities. When looking at specific tasks, it emerges clearly that most of complementarity operates through substitution; in other words, some tasks will be executed by AI and machines while humans will concentrate on others. AI and humans will complement each other in the same occupation while the task allocation within occupation is likely to change.
6 Conclusions

### v3

@@ 3.3 Step 3: Compute the AI Sustitutability
The TEAI index measures exposure to AI but does not specifically address the issue of task substitution or replacement. To deal with this, we leverage information contained in the motivations obtained in the first step.
We synthesize these, generating a consolidated motivation that summarizes the different justifications regarding the AI exposure rate for each occupation-related task. To obtain a quantitative measure of AI’s involvement in each task, we provide this summarized motivation along with the task title and the associated job occupation title to another LLM,prompting it to return the level of AI engagement on a scale from 1 (“no engagement”) to 5 (“complete automation, replacing human execution”) and the motivation regarding the engagement.
Model choice and Results.
In this step we used a single LLM that was larger than those used in the first step. To ensure the complete reproduction of the results, we chose the Qwen 72B model, which, among open-source models, represents a good trade-off between size and performance. The result is a measure of task substitutability that we aggregated at the occupation level using eq (1), obtaining an index of Task Replacement by AI (TRAI).
@@ 4 Human Evaluation of TEAI/TRAI Indexes
To evaluate the outcomes provided by LLMs, we conducted a user study by recruiting 12 evaluators through Prolific palan2018prolific, an online service offering participants for research studies. It has been recognized as a cost-effective platform that provides high-quality data for online research, making it suitable for studies where reliability and data integrity are priorities douglas2023data, and used in several computer-science related user evaluations, see, e.g., de2024large; koppol2021interaction.
Setting the user study.
We sampled occupations from the O*NET database by SOC first-digit, selecting 20% of occupations and choosing two tasks for each occupation: one with the highest and one with the lowest TEAI/TRAI indexes. This resulted in a total of 200 occupations and 400 tasks. The sample was further divided into four chunks of 100 tasks each. Each sample was composed of job occupation title, task title, TEAI value, TEAI natural language motivation, TRAI value, and TRAI natural language motivations. [^9] To avoid potential biases in human judgment, the evaluators were not informed that AI systems had generated the indexes and summaries they reviewed Each chunk was assigned to exactly three participants, whom we asked to consider each task for each job occupation and evaluate the TEAI and TRAI , with a nominal value [^10] I disagree with both; I agree only with TEAI; I agree only with TRAI; I agree with both indicating their degree of agreement with the indices. We aggregate the results, obtaining agreement performance for TEAI and TRAI. It is important to highlight that these results are not intended to represent true accuracy, but a consistency with what our methodology has estimated. Consensus among the participants was also calculated using the equation 4 to assess the extent to which participants agreed on the assignment of evaluation values. Results are shown in Tab. 2.
Results.
Averagely, evaluators tend to agree with (i) TEAI index 75% of the time (48%+27%); (ii) 71% of the times with TRAI (48%+23%), whilst (iii) only 2% of the times they disagree with the indicators. However, the reader may notice that the level of consensus among reviewer is lower than expected. This is largely because the topic at hand — the impact of AI on tasks and professions — is inherently subjective and influenced by human variability, convictions, and expectations, making evaluations highly personal and, at times, open to interpretation. Nonetheless, the results show that our indices are meaningful as they effectively capture the current capabilities of state-of-the-art AI technologies in supporting and replacing tasks within various jobs. As a result, TEAI, TRAI, and the natural language explanations provide a valuable framework for individuals and organizations to develop policies that integrate AI into careers and the workforce.
@@ Results.
Agree with both 35.84 53.76 32 65.94 48 .69 .72 .54 .72 .68
Agree more with TEAI 27.96 19 45.48 17.03 27 .66 .57 .67 .52 .61
Agree more with TRAI 35.13 26.88 20.35 11.23 23 .75 .76 .70 .64 .71
Disagree with both 1.08 0.36 2.17 5.89 2 .50 .27 .41 .28 .36
Table 2: Human evaluation results. Mean of Coherence and Consensus Levels among reviewers by chunk. Each chunk has been assigned to three reviewers. The closer the consensus is to 1, the more the different reviewers agree.
5 Experimental Results and Outcomes on the US Labor Market
Benchmarking evaluation.
First, we compare our AI index with other existing measures in the literature. Figure 2 shows the correlation between the TEAI index and the well-known measure developed by Frey-Osborne-17, the AI exposure index by Felten-Raj-Seamans-21 and by Webb-23, and the offshorability index developed by Acemoglu-Autor-11. The pairwise correlation is always statistically significant at 5%. It is higher for the AIOE index, much lower for the AI Webb and the offshorability index, and negative for the Frey-Osborne index. This means our measure is broadly consistent with existing measures but captures different elements of the relationship between AI and the labor market. The negative correlation with the Frey and Osborne index can be explained by the latter being a measure of exposure to robotics and computerisation, and more focused on routine tasks. At the same time, generative AI is more focused on cognitive/non-routine tasks.
@@ 5 Experimental Results and Outcomes on the US Labor Market
Figure 3: Correlation with different skill intensity measures.
@@ 5.1 AI and Skills
Next, we explore the relationship between our TEAI index and different skills. Figure 3 displays scatterplots comparing the TEAI index with the intensity of different skill types at occupation level derived from Acemoglu-Autor-11. The graph shows the peculiar nature of AI technologies, which are positively correlated with cognitive analytical and interpersonal skills but negatively correlated with routine manual skills and non-routine manual skills that require physical adaptability. Surprisingly, the correlation with cognitive routine skills is only weakly positive, while it is positive for non-routine manual skills requiring interpersonal adaptability. The results of the figure are purely descriptive, therefore we add a more robust analysis by extracting from ONET the detailed skills associated with each occupation. We group the skills into 4 classes: Cognitive, Social, Problem Solving and Management, and Technical skills. We then develop a skill relevance index for each class at the occupation level by weighting each skill according to its level and importance. [^11] As provided by ONET. The skill relevance index is constructed as follows:
@@
where each observation is a SOC occupation ($i$), $TEAI_{i}$ is our measure of AI exposure, $\bm{S}$ is a 4x1 column vector of skill relevance for each of the skill classes described above at the occupation level, $\bm{\beta}$ is a 4x1 column vector of coefficients, $O$ defines occupation dummies and $\epsilon_{i}$ the error term. We saturate the model with more detailed dummies up to the fifth digit; therefore, the results are identified within each group variation. Table 3 shows the results. The TEAI index is positively correlated with cognitive, problem-solving, and managerial skills, but negatively correlated with social skills, as expected. The relationship with technical skills is very weak and does not survive the inclusion of detailed SOC occupation dummies.
@@
Table 3: OLS estimates of TEAI index on measures of skill intensity. Note: Each observation consists of an occupation. OLS regression using TEAI index as the dependent variable. The independent variables are skill intensities. All the regressions include occupation (SOC) fixed effects at 3, 4, and 5 digits. Robust standard errors in parentheses *** p << 0.001, ** p << 0.01, * p << 0.05
@@ 5.2 AI Employment and Wages
Finally, we explore the relationship between TEAI and labor market outcomes. We begin by analyzing the size and characteristics of workers exposed to AI technologies. First, we divide the distribution of TEAI scores into three tertiles representing high, medium, and low AI exposure. We then compute the degree of exposure of the US population using BLS employment data. Finally, we distinguish between occupational groups and by skill groups within each tertile.
@@
Figure 4: Figures 4(a) and 4(b) display the distribution of exposure to the TEAI index across SOC groups and skill intensity levels, respectively, using U.S. BLS employment data (in millions of workers). For both panels, each bar represents a tertile of the TEAI score distribution.
@@
Figures 4(a) and 4(b) show the results. Overall, in 2023, 34% of US employment is highly exposed to AI technologies, while medium and low exposure represent 32% and 34%, respectively. Our findings do not suggest a polarizing effect of AI exposure as found by Frey-Osborne-17; on the contrary, AI seems to have a more balanced impact on the labor market. This is because our indicator can capture recent advances in AI, such as LLMs, which have affected occupational groups such as management, business, administration, and finance, as well as ICT and science, which are intensive in non-routine cognitive tasks. For example, AI technologies are increasingly being used to diagnose diseases, write reports, code, or brainstorm ideas in management and business. On the contrary, previous studies that focus more on the impact of AI on routine tasks find that these tasks and occupations are less exposed to AI. Grouping occupations by skill intensity shows that in the group with high exposure to AI, 88% of employment is in high-skill jobs; in the group with medium exposure, 53% of employment is in medium-skill jobs, while 40% is in high-skill jobs. In the group with the lowest exposure to AI, 67% are medium-skill jobs and 25% low-skill jobs. Overall, AI exposure disproportionately affects high-skill jobs, which are characterized by the competencies most heavily affected by AI technologies. Next, we analyze the relationship between AI exposure and workers’ characteristics. Figure 5 shows that TEAI exposure is higher for workers with a high level of education, particularly graduates and postgraduates. There is a slight increase in exposure with age, although the variation in exposure is limited after age 30. Men are more exposed than women at all ages. Finally, we assess the relationship between AI exposure, employment, and wages. To compute the medium-term effect of AI in a flexible way, allowing for changes during the estimation period, we compute the log change in employment and wages over a 4-year rolling window from 2003 to 2023.121212The last 4Y variation is therefore 2019-2023. Therefore, for each 4Y window, we run the following:
where $\Delta y_{i,j}$ denotes the 4-year change in log employment and log wages in sector $j$ for occupation $i$, $\bm{Z}$ is a column vector of controls, $\delta_{i}$ denote occupation dummies, $\eta_{j}$ sector dummies and $\epsilon_{i,j}$ the error term. To control for possible endogeneity and omitted variable problems, we include as controls the initial level of employment, the initial level of wage, and wage squared. We also include detailed sector (NAICS) and occupation (SOC) fixed effects in the regression and cluster the errors at the sector level. Figure 6 shows that exposure to AI positively correlates with employment and wage growth throughout the entire period. This suggests that AI technologies complement labor and increase productivity, thereby boosting employment and wages in occupations with greater exposure to AI. The presence of detailed controls at the industry and occupation level allows us to control for factors on the production side (changes in output across industries), on the demand side (changes in product demand across industries) and on the labor supply side (changes in employment across industries and occupations) that are unrelated to AI technologies and that could affect wages and employment. Moreover, the focus on a relatively short period isolates our results from long-term trends within industries and occupations. Therefore, the positive relationship between employment and wages and AI exposure should mean that occupations more exposed to AI have stronger employment and wage growth within the occupation and sector. Our results contrast with those obtained by Acemoglu-Autor_Hazell-Restrepo-22, and Webb-23, who find a negative relationship between employment and wages. The potential reconciliation between our findings and theirs lies, on the one hand, in our construction of a different measure of AI exposure that emphasises more recent advances in AI. On the other hand, our analysis focuses on changes over the last 20 years, whereas theirs focuses on changes over several decades.
@@
(a) shows coefficients of regression of education categories on TEAI exposure. Controls include age, sex, occupation(4d), industry(3d), state and year fixed effects. (b) is a binscatter. The x-axis is the average age of workers in an industry-occupation-state observation in the 2022-18 ACS 5-year sample. Biscatter is computed considering education as a covariate. ACS individual weights are used.
@@
This figure plots the effect of AI score on employment and wage growth. Estimates are from eq. 3, with rolling regression coefficients and 95% confidence intervals of 4-year windows, starting in 2003-2007. The dependent variables are annual growth rates of employment and wages. Employment regression includes the log of the initial period of employment. Wage regression includes a log of initial period employment, log initial period wage and log initial period wage squared. Occupation and sector fixed effects included.
Figure 7: Task exposure and task substitutability. Each dot represents an O*NET occupation. Division lines are constructed using median values of the distribution of both indexes. An interactive demo is available at https://terminatoreconomy.com/home
@@ 5.3 Task Exposure or Substitutability?
Whether AI will be an opportunity or a threat in the future will hinge on whether it complements or substitutes human labor. The measure we have developed so far is relatively agnostic in this respect, as we cannot yet disentangle substitutability from complementarity. In other words, a high TEAI exposure measure does not necessarily imply full substitution of labor by technology, which, on the contrary, may rather complement human activities, leading to higher productivity without displacing labor. As described in section 3 the TRAI index is specifically designed to assess task replacement.
What is the relationship between TRAI and TEAI index?
@@
On the one hand, the TRAI index can be considered a sub-product of the TEAI index, as it is constructed using the motivation summaries that LLMs produce in developing the TEAI. On the other hand, it differs because the task description that it analyzes is different (the TEAI uses ONET task description TRAI uses the summary of LLMs descriptions and motivations) and is designed to assess task substitutability, as the prompt asks LLM to specifically assess the possibility of replacing the human in task performance.
@@
Figure 7 displays the two indexes. Both scores have been rescaled on a scale of 0-100 for better readability. Please note that the scaling does not imply an exact percentage. For example, a score of 80 does not mean that there is an 80% exposure or a replacement rate of 80%. Also, while the figure shows different jobs, our replacement and exposure measures are constructed at task level. Thus, even if an occupation has a high level of replacement, it does not mean that AI can completely replace it. Instead, it suggests that AI can perform a large number of tasks in that occupation, either completely or partially.
@@
The figure shows that the two measures are highly correlated, i.e., high values of the TEAI index are generally associated with high values of the replacement index. Grouping occupations improves the explanatory power of the analysis. For example, most ICT and science occupations display a high level of exposure and task substitutability. This means that for these occupations, AI can execute several tasks with a high degree of confidence. Human activity will still be required, and it is likely that these professionals will concentrate on more critical and high-value-added tasks (such as supervision, analysis, developing solutions, etc.), leaving other tasks to AI. Analogously, most teachers are characterized by a high level of exposure by a low level of substitutability: they are likely to work with AI, while AI will replace few tasks. Similarly, several health and care occupations are characterized by low levels of exposure and substitutability with few exceptions, such as “Medical record specialists” whose tasks are highly replaceable by technology.
This provides some insights about the complementarity vs substitutability debate. Complementarity is often interpreted as AI and humans sharing the same activities. When looking at specific tasks it emerges clearly that most of complementarity operates through substitution; in other words some tasks will be executed by AI and machines while humans will concentrate on others. AI and humans will complement each other in the same occupation, while task allocation within occupations is likely to change.
6 Conclusions and Limitations